### v2

# Nonlinear Landau levels in the almost-bosonic anyon gas

Abstract
We consider the quantitative description of a many-particle gas of interacting abelian anyons in the plane, confined in a trapping potential. If the anyons are modeled as bosons with a magnetic flux attachment, and if the total magnetic flux is small compared to the number of particles, then an average-field description becomes appropriate for the low-energy collective state of the gas. Namely, by means of a Hartree–Jastrow ansatz, we derive a two-parameter Chern–Simons–Schrödinger energy functional which extends the well-known Gross–Pitaevskii / nonlinear Schrödinger density functional theory to the magnetic (anyonic) self-interaction. One parameter determines the total number of self-generated magnetic flux units in the system, and the other the effective strength of spin-orbit self-interaction. This latter interaction can be either attractive/focusing or repulsive/defocusing, and depends both on the intrinsic spin-orbit interaction and the relative length scale of the flux profile of the anyons. Densities and energies of ground and excited states are studied analytically and numerically for a wide range of the parameters and align well with a sequence of exact nonlinear Landau levels describing Jackiw–Pi self-dual solitons. With increasing flux, counter-rotating vortices are formed, enhancing the stability of the gas against collapse. Apart from clarifying the relations between various different anyon models that have appeared in the literature, our analysis sheds new light on the many-anyon spectral problem, and also exemplifies a novel supersymmetry-breaking phenomenon.
quantum statistics, anyons, density functional theory, Thomas-Fermi approximation
pacs:
05.30.Pr, 03.75.Lm, 71.15.Mb, 73.43.-f
Planar physics offers exciting possibilities by opening up for intermediate quantum statistics and anyons LeiMyr-77 ; GolMenSha-80 ; GolMenSha-81 ; Wilczek-82a ; Wilczek-82b ; Wu-84a ; GolMenSha-85 , with potentially groundbreaking applications in topological quantum information and computation Kitaev-03 ; Lloyd-02 ; FreKitLarWan-03 . However, despite recent advances in the experimental realization of individual anyons Bartolomei-etal-20 ; Nakamura-etal-20 ; Google-23 ; ReaDas-24 , a major obstacle to large-scale study of anyon physics is our yet limited theoretical understanding of the basic properties of the anyon gas; cf., e.g., the essential role of Gross–Pitaevskii (GP) theory for the study of Bose–Einstein condensation DalGioPitStr-99 ; BaoCai-12 ; PitStr-16 . Namely, the key to unraveling the fundamental relationship between exchange and exclusion statistics for anyons is their enigmatic $N$-body spectral problem CanGir-90 ; CanJoh-94 ; Khare-05 . While $N=2$ is analytically tractable, exhibiting a simple interpolation between the properties (energy, density, etc.) of bosons and fermions LeiMyr-77 ; Wilczek-82b ; Wu-84a , only a small fraction of the spectrum is known analytically for $N\geq 3$ Wu-84b ; Chou-91a ; Chou-91b and the remaining part has been investigated using numerical methods for small systems HatKohWu-91a ; SpoVerZah-91 ; MurLawBraBha-91 ; SpoVerZah-92 . In the early 1990s, the qualitative behavior of the ideal $N$-anyon spectrum in a harmonic trap was studied ChiSen-92 ; LiBhaMur-92 . It showed a highly nontrivial interpolation in terms of angular momentum, and included some quantitative approximations close to bosons and fermions by means of density functionals. Around the same time, it was realized that pointlike but singular (point-interacting) anyons can exist Girvin-etal-90 ; Grundberg-etal-91 ; ManTar-91 ; SenChi-92 ; BorSor-92 ; KimLee-97 , and may manifest a “noninterpolating” spectrum Chou-91a ; MurLawBhaDat-92 . A possible physical interpretation was given in terms of spin-orbit coupling and Pauli supersymmetry Grundberg-etal-91 ; ChoLeeLee-92 ; ComMasOuv-95 ; Mashkevich-96 . Anyon superconductivity was proposed for certain fractions of the statistics parameter relative to fermions Laughlin-88 ; FetHanLau-89 ; CheWilWitHal-89 ; Wilczek-90 , and a Chern–Simons theory of statistics transmutation developed in the context of the fractional quantum Hall effect (FQHE) ZhaHanKiv-89 ; LopFra-91 ; EzaHotIwa-92 , which in a certain self-dual limit admits exact vortex soliton solutions JacWei-90 ; JacPi-90b ; JacPi-90a ; Hagen-91 ; JacPi-91b ; Jackiw-92 ; Dunne-99 ; HorZha-09 . For a recent encyclopedic overview of known properties of the 2D anyon gas, see Lundholm-23 .
In recent times, precise mathematical methods have been developed to investigate the stationary $N$-body ground state LunSol-13b ; Lundholm-16 ; LunQva-20 and to derive effective functional theories of Chern–Simons–Schrödinger (CSS) and Thomas–Fermi (TF) type LunRou-15 ; CorLunRou-17 ; CorLunRou-18 ; CorDubLunRou-19 ; GirRou-21 ; Girardot-21 . In this Letter, we summarize some recent mathematical findings of physical importance for the abelian and almost-bosonic, interacting $N$-anyon problem AtaLunNgu-24 ; AtaGirLun-25 . They include the derivation of an effective CSS model with precise relations for the physical parameters, its stability and supersymmetric properties, as well as the analysis of its emergent spectrum, numerically and analytically studied with the help of Jackiw–Pi vortex solitons or “nonlinear Landau levels” (NLLs).
Microscopic anyon Hamiltonian. For a microscopic description of the anyon gas, we consider the $R$-extended, NN-anyon Hamiltonian, with $\alpha$ flux units per particle, and with spin-orbit coupling parameter $g$:
$H_{N}:=\sum_{j=1}^{N}\left[\left(-i\nabla_{\mathbf{x}_{j}}+\alpha\mathbf{A}^{R}_{j}\right)^{2}+\frac{g}{2}\alpha B^{R}_{j}+V(\mathbf{x}_{j})\right],$ (1)
where $\mathbf{x}_{j}\in\mathbb{R}^{2}$ is the position of the $j$:th particle, $\mathbf{A}^{R}_{j}$ a planar vector potential in Coulomb gauge with magnetic field $B^{R}_{j}:=\operatorname{\mathrm{curl}}\mathbf{A}^{R}_{j}=2R^{-2}\sum_{k\neq j}{\mathds{1}}_{D(\mathbf{x}_{k},R)}(\mathbf{x}_{j})$ (disks of radius $R$ with flux $2\pi$ at $\mathbf{x}_{k}$), and $V$ is a regular one-body trapping potential, i.e. $V(\mathbf{x})\to\infty$ as $\mathbf{x}\to\infty$. Here we will consider the harmonic trap $V_{\mathrm{h}}(\mathbf{x}):=\frac{1}{2}m\omega^{2}|\mathbf{x}|^{2}$. We normalize units $\hbar=2m=\frac{\omega}{2}=1$, and allow pointlike anyons by taking $R\to 0$ and ${\mathds{1}}_{D(\mathbf{x}_{k},R)}/(\pi R^{2})\to\delta_{\mathbf{x}_{k}}$, the Dirac measure. Note that, even in the pointlike case at fixed $N$, there are different anyon types specified by $g$, where $g=\pm 2$ defines “ideal anyons” (regular/hard) and “kreinyons” (singular/soft), respectively; see Lundholm-23 ; AtaGirLun-25 for review. The Hamiltonian (1) acts on symmetric wave functions $\Psi_{N}\in L^{2}_{\mathrm{sym}}(\mathbb{R}^{2N})$ s.t. $\alpha=0$ corresponds to ideal bosons and $\alpha=1$, $R=0$ to ideal spinless fermions.
Mesoscopic CSS average-field functional. If $N\gg 1$ while the total flux $\alpha N\simeq\beta$ is of order 1, then on intermediate length scales $R\ll\ell\ll\omega^{-1/2}$ the interacting anyon gas is well described by the following Chern–Simons–Schrödinger (CSS) energy functional:
$\mathcal{E}_{\beta,\gamma,V}[\phi]:=\int_{\mathbb{R}^{2}}\left[\bigl|(-i\nabla+\beta\mathbf{A}_{\varrho})\phi\bigr|^{2}+\gamma|\phi|^{4}+V|\phi|^{2}\right].$ (2)
Here $\gamma$ is an effective coupling parameter, $B_{\varrho}=\operatorname{\mathrm{curl}}\mathbf{A}_{\varrho}=2\pi\varrho$ is a self-consistently generated magnetic field with density distribution $\varrho=|\phi|^{2}$, and $\phi\in L^{2}(\mathbb{R}^{2})$ a one-body collective state, the CSS wave function (see Fig. 1). We denote the ground-state energy of (2), i.e. its infimum over regular $\phi$ subject to the normalization constraint $\left\lVert\phi\right\rVert=1$, by $E(\beta,\gamma,V)$. For $\beta=0$, (2) reduces to the GP energy functional in zero magnetic field.
Figure 1: Density and phase of an approximate minimizer for $E(\beta=10,\gamma=20\pi,V_{\mathrm{h}})\approx 9.066$.
To approximate the low-energy stationary states of the anyon gas, we use a Hartree–Jastrow ansatz for (1):
$\Psi_{N}:=C_{N}\prod_{i<j}f_{R/\ell}(|\mathbf{x}_{i}-\mathbf{x}_{j}|/\ell)\prod_{k=1}^{N}\phi(\mathbf{x}_{k}),$ (3)
where $C_{N}>0$ is a normalization constant s.t. $\left\lVert\Psi_{N}\right\rVert=1$, $\phi\in L^{2}(\mathbb{R}^{2})$ a smooth variational one-body state, and
$f_{s}(r):=\begin{cases}\frac{4s^{\alpha}}{2\left(1+s^{2\alpha}\right)+g\left(1-s^{2\alpha}\right)},\quad&\textup{if }0\leq r\leq s,\\ \frac{(2+g)r^{\alpha}+(2-g)s^{2\alpha}r^{-\alpha}}{2\left(1+s^{2\alpha}\right)+g\left(1-s^{2\alpha}\right)},\quad&\textup{if }s\leq r\leq 1,\\ 1,\quad&\textup{if }r\geq 1.\end{cases}$
In AtaGirLun-25 it is shown that, if $\alpha\simeq\beta/N$, $0\leq R=\mathcal{O}(e^{-\sigma N})\ll\ell=\mathcal{O}(N^{-\tau})$, and $\beta>0$, $\sigma\geq 0$, $\tau>2$, $g\geq 0$, then the energy per particle
$\lim_{N\to\infty}N^{-1}\langle\Psi_{N}|H_{N}|\Psi_{N}\rangle=\mathcal{E}_{\beta,\gamma,V}[\phi],$ (4)
where
$\gamma=2\pi\beta\frac{1+\frac{g}{2}-(1-\frac{g}{2})e^{-2\beta\sigma}}{1+\frac{g}{2}+(1-\frac{g}{2})e^{-2\beta\sigma}}.$ (5)
Note that, in the “supersymmetric” case $g=\pm 2$ this yields exactly $\gamma=\pm 2\pi\beta$ for all relative scales $\sigma\sim N^{-1}\log\frac{\ell}{R}$. In the relatively dense limit $\sigma\to 0$ it becomes $\gamma=g\pi\beta$ whereas in the dilute limit $\sigma\to+\infty$, g>2g>-2 it is $\gamma=2\pi\beta$. We can also notice that in the non-interacting/“spinless” limit $g\to 0$ it is $\gamma=2\pi\beta\tanh(\beta\sigma)$, and $\gamma=2\pi\beta\coth(\beta\sigma)$ in the case of hard disks $g\to+\infty$, $\sigma>0$. This extends the previously studied mean-field/Hartree case $\sigma=0=g$, $f_{s}=1$, for which matching lower bounds are also available LunRou-15 ; Girardot-21 ; Visconti-25 .
[…]
Figure 2: Stability holds if and only if $E(\beta,\gamma,V)>$-$\infty$. NLLs exist at $\beta\in 2\mathbb{Z}\setminus\{0\}$ with the critical coupling $\gamma=-2\pi|\beta|$.
The derivation of the precise formula (5) was motivated by a previous result by Mashkevich Mashkevich-96 concerning the asymptotic 2-body energy. Namely, we define a relative s-wave “scattering energy” $E_{2}$ as the infimum of
$\int_{D(0,1)}\left[|\nabla f|^{2}+\alpha^{2}\frac{|\mathbf{x}|^{2}}{\max\{R,|\mathbf{x}|\}^{4}}|f|^{2}+\alpha g\frac{{\mathds{1}}_{D(0,R)}}{R^{2}}|f|^{2}\right],$
with $f\in L^{2}$ regular on the unit disk with $f|_{r=1}=1$. Then our function $f_{s}$ captures this scattering energy to order $\alpha^{2}$, and its contribution to the $N$-body energy per particle as $N\to\infty$ (with $\binom{N}{2}$ pairs and relative mass 2) is $(N-1)E_{2}\simeq\gamma$ as in (5); see AtaGirLun-25 . Compare to the scattering length and energy for dilute Bose gases LieSeiSolYng-05 ; Solovej-25 , and note that the anyonic vector interaction is long ranged but suppressed by small $\alpha$, while the spin-orbit scalar interaction is relatively short ranged.
Stability and supersymmetry. For each $\beta$ there is a critical attractive coupling $\gamma$ for the CSS model beyond which the gas collapses; we denote it by $\gamma_{*}(\beta)$. Thus, for any $(\beta,\gamma)$ s.t. $\gamma\geq-\gamma_{*}(\beta)$, we have $E(\beta,\gamma,V)>$-$\infty$ (stability), while if $\gamma<-\gamma_{*}(\beta)$ then $E(\beta,\gamma,V)=$-$\infty$ (instability); see Fig. 2. By scaling, it is sufficient to consider $V=0$ and $E(\beta,\gamma,0)=0$ resp. -$\infty$, and the critical coupling $\gamma_{*}(\beta)$ is exactly the smallest (by infimum) ratio between the magnetic-kinetic self-energy $\mathcal{E}_{\beta,0,0}[\phi]$ and the scalar/electrostatic self-energy $\int|\phi|^{4}$ among all normalized and regular $\phi$. Note the complex conjugation symmetry $\phi\to\overline{\phi}$, $\beta\to-\beta$, implying $\gamma_{*}(-\beta)=\gamma_{*}(\beta)$.
$\beta$
$E(\beta,\gamma=-2\pi\beta,V_{\mathrm{h}})$
$\beta$
$E(\beta,\gamma=-0.9\times 2\pi\beta,V_{\mathrm{h}})$
Figure 3: Analytical and numerical energies for attractive gas $\gamma=2\pi\beta\kappa<0$ at $\kappa=-1$ (“kreinyons”) resp. $\kappa=-0.9$.
$\beta$
$E(\beta,\gamma=0,V_{\mathrm{h}})$
$\beta$
$E(\beta,\gamma=2\pi\beta,V_{\mathrm{h}})$
NLL degree:
1
2
3
4
5
# vortices:
0
1
2
3
4
5
6
8
9
Figure 4: Analytical and numerical energies for repulsive gas $\gamma=2\pi\beta\kappa\geq 0$ at $\kappa=0$ resp. 1 (ideal anyons). See the text.
At $\beta=0$, the critical attraction for stability of the GP functional is the constant $\gamma_{*}(0)=:C_{\rm LGN}\approx 0.931\times 2\pi$ (see Weinstein-83 ), realized by “Townes’ soliton” ChiGarTow-64 ; Fibich-15 , the real and unique solution (mod scaling and translation) at $E(0,-C_{\rm LGN},0)=0$. In AtaLunNgu-24 it was shown that $\gamma_{*}(\beta)=2\pi\beta$ for $\beta\geq 2$, while for intermediate $0<\beta<2$ it holds that $\gamma_{*}(\beta)>\max\{C_{\rm LGN},2\pi\beta\}$. Further, although as a limit $E(\beta,-2\pi\beta,0)=0$ for all $\beta\in\mathbb{R}$, zero-energy solutions only exist for $\beta=2,4,6,\ldots$, and we refer to these as “nonlinear Landau levels” (NLLs). Namely, at the self-dual coupling $\gamma=-2\pi\beta$, with $B=2\pi\beta\varrho$,
$\mathcal{E}_{\beta,-2\pi\beta,0}[\phi]=\bigl\langle\phi\big|(-i\nabla+\mathbf{A})^{2}-B\big|\phi\bigr\rangle.$ (6)
If $\mathbf{A}=\frac{1}{2}(\partial_{2},-\partial_{1})\psi$, where the superpotential $\psi$ solves $\Delta\psi=2B$, we can use the supersymmetric factorization
$\mathcal{E}_{\beta,-2\pi\beta,0}[\phi]=\int_{\mathbb{R}^{2}}\bigl|(\partial_{1}-i\partial_{2})(e^{-\psi/2}\phi)\bigr|^{2}e^{\psi}.$ (7)
Cf. the standard, linear Landau problem, where $B>0$ is a constant and where the zero-energy subspace, the lowest Landau level, consists of all normalizable functions on the form $\phi=\overline{f(z)}e^{\psi/2}$, where $f$ is analytic in $z=x_{1}+ix_{2}\in\mathbb{C}$, and $\psi(\mathbf{x})=-B\lvert\mathbf{x}\rvert^{2}$. In our problem, $\psi$ in (7) solves the nonlinear and nonlocal, generalized Liouville equation $\Delta\psi=4\pi\beta|f|^{2}e^{\psi}$, yielding the exact solutions AtaLunNgu-24
$f\propto P^{\prime}Q-PQ^{\prime},\qquad\phi_{P,Q}:=\sqrt{\frac{2}{\pi\beta}}\,\frac{\overline{P^{\prime}Q-PQ^{\prime}}}{|P|^{2}+|Q|^{2}},$ (8)
where $P,Q$ are linearly independent and coprime complex polynomials with $n:=\max(\deg P,\deg Q)=\beta/2$. This complete set of regular minimizers includes all the self-dual soliton solutions found by Jackiw and Pi: initially, radial vortex rings with $P=z^{n}$ and $Q=1$ JacPi-90b , and eventually, the more general rationals $P/Q$ JacPi-90a . Thus, along the critical line $\gamma=-2\pi\beta$, $\beta\in\mathbb{R}$, the sequence of points $\beta=2n$ marks a phase transition from strictly positive (broken supersymmetry) to zero energy and increasing degree of supersymmetry, exhibited by NLL manifolds of real dimension $4n$; compare Figs. 2 and 3 (top).
Macroscopic energy and density in a harmonic trap: numerics. For $\beta\gg 1$ and $\gamma\propto\beta$ in the stable regime, to minimize the energy we anticipate the formation of a locally homogeneous vortex lattice on scales $\ell\sim\beta^{-1/4}$; cf. CorLunRou-17 ; CorLunRou-18 ; CorDubLunRou-19 . In contrast to Abrikosov lattices in strong external magnetic fields AftBlaDal-05 ; Fetter-09 , the lattice spacing here follows the density profile sharply. Therefore, by local density approximation on the macroscopic scale of the trap, we expect an approximate Thomas–Fermi (TF) description to hold for a coarse-grained average density $\tilde{\varrho}$:
$\mathcal{E}_{\beta,\gamma,V}[\phi]\ \approx\ \mathcal{E}^{\mathrm{TF}}_{G,V}[\tilde{\varrho}]:=\int_{\mathbb{R}^{2}}\left[G{\tilde{\varrho}}^{2}+V\tilde{\varrho}\right$], (9)
subject to $\int\tilde{\varrho}=1$, $V=V_{\mathrm{h}}$, and a constant $G$ depending on both $\beta$ and $\gamma$. The minimum energy $E(\beta,\gamma,V_{\mathrm{h}})\simeq E^{\mathrm{TF}}(G)=\frac{4}{3}\sqrt{G/\pi}$ and the minimizer $\varrho\approx\tilde{\varrho}^{\mathrm{TF}}=(2G)^{-1}(\lambda^{\rm TF}-|\mathbf{x}|^{2})_{+}$. For comparison, $G=2\pi N$ yields the asymptotic density and energy per particle of $N$ spinless fermions in the harmonic trap.
In Figs. 34 we study the energy spectrum of (2) along the line $\gamma=2\pi\beta\kappa$, for $\beta\in[0,10]$ and four different slopes $\kappa$. When $\beta=\gamma=0$ it is simply the Schrödinger energy of the 2D harmonic oscillator, with spectrum $2,4,6,\ldots$. As upper bounds we have computed the analytical energies of the first few radial vortex rings ($n=2,3,4,5$; optimally rescaled) and of an approximation to the “versiera” at $n=1$ with finite potential energy; these are the solid curves in the plots; see Supplementary Material (SM) for details. We have compared to the ground (Gaussian, at $n=1$) and first excited (single counter-rotating vortex, at $n=2$) states of the harmonic oscillator (again optimally rescaled; dashed curves). Further, general energies have been computed numerically (solid markers) by means of a gradient descent method, which is an extension of that used in CorDubLunRou-19 to the scalar self-interaction, with $(\beta,\gamma)$ in the stable regime. We have tested its accuracy with $\beta=0$ and various $\gamma$ to reproduce GP energies; cf. (BaoCai-12, , Table 3.5). Also, at $\gamma=0$ we reproduce the minimal energies of CorDubLunRou-19 , i.e. the minima follow closely the estimate $E^{\mathrm{TF}}(2\pi c\beta)$ where the numerical fit $c=2\sqrt{\pi}/3\approx 1.18$ is a renormalizing “Abrikosov factor” originating from the small-scale variations of the emerging vortex lattice. Interestingly, this behavior holds also for small $\beta$ all the way down to $\beta\gtrsim 2$. The critical $\beta$ for nucleation of a single vortex is just slightly below 3.
For a general estimate of the ground-state energy at $\gamma=2\pi\beta\kappa$, $\kappa\geq-1$, we have compared with the linear interpolation $E^{\mathrm{TF}}(2\pi c(1+\kappa)\beta)=E^{\mathrm{TF}}(2\pi\beta)\sqrt{c}\sqrt{1+\kappa}$ (dash-dotted curve), which provides a fit to numerics to within a few percent for $\beta\gtrsim 3$ and $0.9\lesssim\kappa\lesssim 2$. The estimate given by ChiSen-92 for $\kappa=1$ is ca 7% smaller. An absolute lower bound for the energy (assuming regularity and singlevaluedness) is provided by $\gamma_{*}(\beta)\to 2\pi\beta$ and thus differs by a factor $\sqrt{c}$, i.e. 9% (dotted curve; which for ideal anyons is still 41% higher than the ideal Fermi energy). We also conjecture that the real-valued minimizer at $\kappa=\pm 1$ has a linearly interpolating energy $|2\pm\beta|$. This, as well as other linear features of the $\kappa=\pm 1$ spectra, resembles the known linear/noninterpolating exact states of the $N$-anyon spectrum Chou-91a ; MurLawBhaDat-92 ; ChiSen-92 ; AtaGirLun-25 .
For $\beta=2n\in\{4,6,8\}$ we have used analytical expressions for the NLLs (from solving the inverse Wronskian problem $P,Q\mapsto W(P,Q):=P^{\prime}Q-PQ^{\prime}$) and numerical integration together with optimization of the placement of zeros/vortices to estimate their minimal energy
$\mathcal{E}_{\beta,\gamma,V_{\mathrm{h}}}[\phi_{P,Q}]=\int_{\mathbb{R}^{2}}\left[(4\pi n+\gamma)|\phi_{P,Q}|^{4}+V_{\mathrm{h}}|\phi_{P,Q}|^{2}\right].$ (10)
These are the hollow markers in the plots (again, see SM for details). Although for small $\beta$ it is very costly to break the radial symmetry, there is a trend that for larger $\beta$ and smaller $\gamma$ this cost can be reduced in sufficiently symmetric vortex configurations, and indeed for $\beta=8$, $\kappa=-0.9$ an NLL state with 4 separated vortices has lower energy than any of the radial vortex rings.
Conclusions and outlooks. We have derived and analyzed an effective CSS model of the interacting many-anyon gas in the bosonic limit, which manifests a spectrum of energies/states that are well describable using a family of exact NLLs (8). Although we have here focused on small flux $\beta\lesssim 10$, the Matlab part of our numerics extends to ca $\beta\lesssim 140$ (see Table 12 in SM). Our study of exact NLLs is mainly limited by the increasing complexity of the inverse Wronskian problem, however we may extend relatively easily to some symmetry-reduced cases for $\beta\geq 10$. Ongoing work also includes external magnetic fields and the Abrikosov problem for self-generated vortex lattices. Further, in our analysis of NLLs we have assumed regularity and singlevaluedness of CSS wave functions $\phi$. We believe that in order to obtain even lower energies, to eventually match the ideal Fermi energy at $\alpha\to 1$, we must relax these assumptions. This also leads to the study of systems of CSS functions and Liouville–Toda-type equations (cf. DunJacPiTru-91 ; Dunne-92 ; Kim-etal-93 ), which will be addressed in future work. Some new insight into the fermionic side of the problem, via fermionic Hartree and magnetic TF theory, has been recently discussed in LevLunRou-25 (see also Hu-etal-21 ).
In conclusion, we believe that important keys to understanding the complex interpolation between the almost-bosonic and almost-fermionic anyon spectra will come via the NLLs and their possible extension to more general functions. Also, exploring the two-parameter energy landscape may uncover fruitful and highly nontrivial relations between the CSS, GP, TF, and other models.
Acknowledgments.
We are grateful to Romain Duboscq for the original Matlab code upon which much of our numerical work is based. A.A., A.E., F.G. and D.L. thank the Department of Mathematics of Politecnico di Milano for its kind hospitality during the spring of 2025 intensive period “Quantum Mathematics at Polimi”. D.L. also thanks the GSSI L’Aquila for its hospitality. All of us have benefited from discussions at the Institut Mittag-Leffler workshop “Anyons from small to large scales” in July 2025. Financial support from the Swedish Research Council (D.L., grant no. 2021-05328, “Mathematics of anyons and intermediate quantum statistics”) is gratefully acknowledged. A.A. thanks Kaj Nyström for the financial support for traveling to the relevant conferences. The author T.G. acknowledges the Gran Sasso Science Institute for the financial support making this work possible. A.E. gratefully thanks the Center for Interdisciplinary Mathematics at Uppsala University as well as C.F. Liljewalchs stipendiestiftelse for their financial support. D.-T.N. was funded by the Postdoctoral Scholarship Programme of Vingroup Innovation Foundation (VINIF), VinUniversity, code VINIF.2025.STS.15, and also gratefully acknowledges the support of the Knut and Alice Wallenberg Foundation through the grant of Prof. Andreas Strömbergsson during his postdoctoral period at Uppsala University (2022–2024).
Supplementary material
Here we give all the information necessary to reproduce the plots of Figs. 34. After setting our conventions, we first derive some explicit analytical NLL states (8) by solving the inverse Wronskian problem. Then we analytically compute the energies of some related trial states, including a reference Gaussian, the first excited harmonic oscillator energy eigenstates, powers of the “versiera” ansatz, as well as NLL “vortex rings”. Finally, we give tables of numerical data.
Notation and choice of gauge
The $R$-extended anyon gauge vector potential is defined by $\mathbf{A}^{R}_{j}:=\sum_{k\neq j}\frac{(\mathbf{x}_{j}-\mathbf{x}_{k})^{\perp}}{|\mathbf{x}_{j}-\mathbf{x}_{k}|_{R}^{2}}$ where $\mathbf{x}^{\perp}:=(-x_{2},x_{1})$ denotes the counterclockwise $90^{\circ}$ rotation for $\mathbf{x}=(x_{1},x_{2})\in\mathbb{R}^{2}$, and $|\mathbf{x}|_{R}:=\max\{R,|\mathbf{x}|\}$. The effective gauge potential for a given density distribution $\varrho\in L^{1}(\mathbb{R}^{2};\mathbb{R}_{+})$ is defined by the convolution
$\mathbf{A}_{\varrho}(\mathbf{x}):=\int_{\mathbb{R}^{2}}\frac{(\mathbf{x}-\mathbf{y})^{\perp}}{|\mathbf{x}-\mathbf{y}|^{2}}\varrho(\mathbf{y})\,d\mathbf{y}.$
In both vector potentials we have chosen the Coulomb gauge $\nabla\cdot\mathbf{A}=0$, and they generate the magnetic fields
$B^{R}_{j}=\operatorname{\mathrm{curl}}_{\mathbf{x}_{j}}\mathbf{A}^{R}_{j}=\begin{cases}\frac{2}{R^{2}}\sum_{k\neq j}{\mathds{1}}_{D(\mathbf{x}_{k},R)}(\mathbf{x}_{j}),&R>0,\\ 2\pi\sum_{k\neq j}\delta(\mathbf{x}_{j}-\mathbf{x}_{k}),&R=0,\end{cases}$
respectively
$B_{\varrho}=\operatorname{\mathrm{curl}}\mathbf{A}_{\varrho}=2\pi\varrho$
(in two dimensions, the magnetic field is a (pseudo-) scalar quantity), where we used that $\Delta\log|\mathbf{x}|=2\pi\delta(\mathbf{x})$. Throughout, we use the notation $D(\mathbf{x},r)$ for a disk of radius $r$ at $\mathbf{x}$, and ${\mathds{1}}_{\Omega}$ for the characteristic function or projection operator supported on $\Omega$. Also, $x_{+}:=\max\{0,x\}$. Sometimes we use complex notation for the coordinates $z=x_{1}+ix_{2}=re^{i\theta}\in\mathbb{C}$, $\bar{z}=x_{1}-ix_{2}$.
Note that, because of our gauge choice and vanishing condition $\lim_{\mathbf{x}\to\infty}\mathbf{A}_{\varrho}=0$, for a radially symmetric density $\varrho(\mathbf{x})=\varrho(r)$, we must have
$\mathbf{A}_{\varrho}(\mathbf{x})=A(r)\mathbf{e}_{\theta},$ (11)
for some radial function $A(r)$. Stokes’ theorem then implies that
$2\pi rA(r)=\oint_{\partial D(0,r)}\mathbf{A}_{\varrho}\cdot d\mathbf{s}=\int\limits_{D(0,r)}\operatorname{\mathrm{curl}}\mathbf{A}_{\varrho}=2\pi\int\limits_{D(0,r)}\varrho,$
so
$A(r)=\frac{1}{r}\int\limits_{D(0,r)}\varrho=\frac{2\pi}{r}\int_{0}^{r}\varrho(s)s\,ds.$ (12)
The Euler–Lagrange equation for a minimizing ground state or a stationary point $\phi$ of (2) is the CSS equation
$\left[(-i\nabla+\beta\mathbf{A}_{\varrho})^{2}-2\beta\nabla^{\perp}(\log|\cdot|)*\mathcal{J}_{\phi}+2\gamma\varrho+V\right]\phi=\mu\phi,$ (13)
where
$\mathcal{J}_{\phi}:=\frac{1}{2}\overline{\phi}(-i\nabla+\beta\mathbf{A}_{\varrho})\phi+\text{h.c.}$
and
$\mu=2\mathcal{E}_{\beta,\gamma,V}[\phi]-\int_{\mathbb{R}^{2}}\left[|\nabla\phi|^{2}+V\varrho-\beta^{2}\left|\mathbf{A}_{\varrho}\right|^{2}\varrho\right];$ (14)
see (CorLunRou-17, , Appendix) and (AtaLunNgu-24, , Remark 3.11).
.1 NLL states and inverse Wronskian problem
The general formula (8) for NLL states is parametrized by the space of solutions $(P,Q)$ of the inverse Wronskian problem
$W(P,Q):=P^{\prime}Q-PQ^{\prime}=f,$ (15)
where $f$ is a given polynomial and $P$ and $Q$ are coprime (and linearly independent) polynomials satisfying the degree condition $\max\{\deg P,\deg Q\}=n=\beta/2>0$.
Due to the determinantal property of the Wronskian and the symmetry of the amplitude $|P|^{2}+|Q|^{2}$, the corresponding NLL state $\phi_{P,Q}$ given by (8) is invariant under pointwise action of the group $\mathrm{SU}(2)\times\mathbb{R}^{+}$ on the space of polynomial pairs $(P,Q)$. It follows that we may without loss of generality assume that either $P$ or $Q$ is monic, and, e.g., that $\deg Q<\deg P=n=\beta/2$. We may also simplify the problem (15) by adding any multiple of $Q$ to $P$. If we identify states that differ only by multiplication by a constant phase factor $e^{i\theta}$, we have a further $\mathrm{U}(1)$-symmetry, for a full $\mathrm{U}(2)\times\mathbb{R}^{+}$-symmetry in total. See AtaLunNgu-24 for further details.
Thus, given the location of vortices/zeros $w_{k}\in\mathbb{C}$, i.e. $f(z)\propto\prod_{k}(z-w_{k})$, we may enumerate all NLL solutions with these vortices by
$(\tilde{P},\tilde{Q})=(P+\eta Q,\lambda^{-1}Q),\qquad\eta\in\mathbb{C},\quad\lambda>0,$
and
$\phi_{\tilde{P},\tilde{Q}}=\frac{1}{\sqrt{\pi n}}\frac{\overline{W(P,Q)}}{\lambda\lvert P+\eta Q\rvert^{2}+\lambda^{-1}\lvert Q\rvert^{2}},$ (16)
where $(P,Q)$ are particular solutions to the Wronskian problem (15) of the simplified form.
Notice that the number of vortices for a given NLL state $\phi_{P,Q}$ (counting multiplicities) is precisely $\deg W(P,Q)$, which for fixed $\beta/2=n=\max\{\deg P,\deg Q\}$ is restricted by the inequalities
$n-1\leq\deg W(P,Q)\leq 2n-2.$ (17)
Hence the $n$:th NLL space contains states with between n1n-1 and 2n22n-2 vortices. We will now list some known NLL solutions for various $n$.
For any positive integer $n=\beta/2=1,2,3,\ldots$, we may consider the degenerate case
$f(z)=n(z-w)^{n-1},\qquad w\in\mathbb{C}.$
A particular solution to the inverse Wronskian problem $W(P,Q)=f$ is then
$P(z)=(z-w)^{n},\qquad Q(z)=1.$ (18)
The corresponding state on the form (16) is the “vortex ring
$\phi_{\tilde{P},\tilde{Q}}(z)=\sqrt{\frac{n}{\pi}}\frac{(\bar{z}-\bar{w})^{n-1}}{\lambda\lvert(z-w)^{n}+\eta\rvert^{2}+\lambda^{-1}},$ (19)
for $\lambda>0$ and $\eta\in\mathbb{C}$. For $\eta=0$, the corresponding density is radially symmetric about $w$. These solutions (for $w=\eta=0$) were first described by Jackiw and Pi in JacPi-90b . For $n>1$, they have a vortex (zero) of degree n1n-1 at $z=w$, but in the special case $n=1$, we get the real-valued and positive “versiera” solution
$\phi(z)=\frac{1}{\sqrt{\pi}}\frac{1}{\lambda\lvert z-w\rvert^{2}+\lambda^{-1}},$ (20)
which lacks vortices (note that the parameter $\eta$ can be absorbed into $w$ in this case). The family of versiera solutions, parametrized by the scale $\lambda>0$, center $w\in\mathbb{C}$, and a constant phase, is the complete manifold of NLL states for $n=1$.
For fixed $n=\beta/2$, the $4n$-dimensional manifold of NLL states foliates over the degree $d=\deg f$, $n-1\leq d\leq 2(n-1)$. For $n=1$, we must have $d=0$, so the versiera solutions are the only possible solutions. For $n=2$ and $d=1$, we have the particular solution
$f(z)=z-w,\quad P(z)=z^{2}-2wz,\quad Q(z)=\frac{1}{2},$ (21)
corresponding to the family of vortex ring solutions (19) of degree 1, and for $d=2$ we have the particular solution
$f(z)$ $=(z-w_{1})(z-w_{2}),$
$P(z)$ $=z^{2}-w_{1}w_{2},$ (22)
$Q(z)$ $=z-\frac{w_{1}+w_{2}}{2}.$
We must have $w_{1}\neq w_{2}$, for otherwise $P$ and $Q$ are not coprime. Hence, these states have two separated vortices located at $w_{1}$ and $w_{2}$. With the additional degrees of freedom in (16), these two families of solutions exhaust the NLL manifold for $n=2$ (this was discussed in (AtaLunNgu-24, , Section 2.3)).
For $n=3$, we have $2\leq d\leq 4$ by (17). For $d=2$, we have the particular solution
$f(z)=(z-w_{1})(z-w_{2}),$
$P(z)=z^{3}-3\frac{w_{1}+w_{2}}{2}z^{2}+3w_{1}w_{2}z,$ $Q(z)=\frac{1}{3},$ (23)
with $w_{1},w_{2}\in\mathbb{C}$ arbitrary. If $w_{1}=w_{2}$, this is a vortex ring (19) of degree 2. For $d=3$, we have two particular solutions:
$f(z)$ $=(z-w_{1})(z-w_{2})(z-w_{3}),$
$P(z)$ $=z^{3}-A_{\pm}z^{2}+2s_{3},$ (24)
$Q(z)$ $=\frac{1}{2}z-\frac{1}{6}A_{\mp},$
where
$A_{\pm}$ $:=s_{1}\pm\sqrt{s_{1}^{2}-3s_{2}}.$
Here
$s_{k}=s_{k}(w_{1},\ldots,w_{d}):=\sum_{1\leq i_{1}<\cdots<i_{k}\leq d}w_{i_{1}}\cdots w_{i_{k}}$
is the $k$:th elementary symmetric polynomial in the variables $w_{1},\ldots,w_{d}$. In particular,
$s_{1}(w_{1},w_{2},w_{3})$ $=w_{1}+w_{2}+w_{3},$
$s_{2}(w_{1},w_{2},w_{3})$ $=w_{1}w_{2}+w_{1}w_{3}+w_{2}w_{3},$
$s_{3}(w_{1},w_{2},w_{3})$ $=w_{1}w_{2}w_{3}.$
For $d=4$, we have the two particular solutions
$f(z)$ $=(z-w_{1})(z-w_{2})(z-w_{3})(z-w_{4}),$
$P(z)$ $=z^{3}-\frac{1}{2}B_{\mp}z+\frac{1}{2}s_{3},$ (25)
$Q(z)$ $=z^{2}-\frac{1}{2}s_{1}z+\frac{1}{6}B_{\pm},$
where
$B_{\pm}$ $:=s_{2}\pm\sqrt{s_{2}^{2}-3s_{1}s_{3}+12s_{4}}.$
The above particular solutions yield all NLL states for $n=3$ via (16), and $P,Q$ are coprime at least if all the roots $w_{j}$ are disjoint (consider $f^{\prime}=P^{\prime\prime}Q-PQ^{\prime\prime}$).
For $n=4$, we will be content with giving some special solutions with additional symmetries that have been used in the numerics. With $d=\deg f$ as usual, we have $3\leq d\leq 6$. For $d=3,4$, we can for example place the vortices at a distance $R$ from the origin along the roots of unity, $w_{k}=Re^{i2\pi k/d}$. This yields the particular solutions
$f(z)=z^{3}-R^{3},\quad P(z)=z^{4}-4R^{3}z,\quad Q(z)=\frac{1}{4},$ (26)
for $d=3$, and
$f(z)=z^{4}-R^{4},\quad P(z)=z^{4}+3R^{4},\quad Q(z)=\frac{1}{3}z,$ (27)
for $d=4$. For $d=5$, we may place one vortex at the origin and the other four along the roots of unity, yielding the particular solution
$f(z)=z(z^{4}-R^{4}),\quad P(z)=z^{4}+R^{4},\quad Q(z)=\frac{1}{2}z^{2}.$ (28)
For $d=3$ we can allow $R\geq 0$, while for $d=4,5$ we require $R>0$ for coprimality.
Analytical energies of trial states
We will now calculate the exact energy $\mathcal{E}[\phi]=\mathcal{E}_{\beta,\gamma,V}[\phi]$ analytically for certain trial CSS states $\phi$ in the harmonic trap $V(z)=V_{\mathrm{h}}(z)=\lvert z\rvert^{2}$. The parameters $\beta$ and $\gamma$ are considered fixed.
First we make some general observations. By expanding the kinetic-magnetic term $\lvert(\nabla+i\beta\mathbf{A}_{\varrho})\phi\rvert^{2}$, the energy functional can be recast into the form
$\mathcal{E}[\phi]=\int_{\mathbb{R}^{2}}\left[\lvert\nabla\phi\rvert^{2}+2\beta\mathbf{A}_{\varrho}\cdot\mathbf{J}_{\phi}+\beta^{2}\lvert\mathbf{A}_{\varrho}\rvert^{2}\varrho+\gamma\varrho^{2}+V\varrho\right],$ (29)
where
$\mathbf{J}_{\phi}:=\frac{i}{2}\left(\phi\nabla\overline{\phi}-\overline{\phi}\nabla\phi\right)$ (30)
is the probability current induced by $\phi$. When $\phi$ has a constant phase, $\mathbf{J}_{\phi}=0$.
Now we turn to the question of finding the optimal length scale in the harmonic potential $V_{\mathrm{h}}$, or equivalently the optimal scale of the trial state. If $\phi$ is a normalized trial state and $\lambda>0$, let
$\phi_{\lambda}(z):=\lambda\phi(\lambda z).$ (31)
The rescaling $\phi\to\phi_{\lambda}$ induces rescalings of the quantities derived from $\phi$:
$\nabla\phi_{\lambda}(z)$ $=\lambda^{2}(\nabla\phi)(\lambda z);$
$\varrho_{\lambda}(z)$ $:=\lvert\phi_{\lambda}(z)\rvert^{2}=\lambda^{2}\varrho(\lambda z);$
$\mathbf{A}_{\lambda}(z)$ $:=\mathbf{A}_{\varrho_{\lambda}}(z)=\lambda\mathbf{A}_{\varrho}(\lambda z);$
$\mathbf{J}_{\lambda}(z)$ $:=\mathbf{J}_{\phi_{\lambda}}(z)=\lambda^{3}\mathbf{J}_{\phi}(\lambda z).$
As a result, if we let
$K_{\lambda}$ $:=\int_{\mathbb{R}^{2}}\left[\lvert(\nabla+i\beta\mathbf{A}_{\lambda})\phi_{\lambda}\rvert^{2}+\gamma\varrho_{\lambda}^{2}\right];$ (32)
$U_{\lambda}$ $:=\int_{\mathbb{R}^{2}}V_{\mathrm{h}}\varrho_{\lambda},$ (33)
then $K_{\lambda}=\lambda^{2}K_{1}$ and $U_{\lambda}=\lambda^{-2}U_{1}$, while $\int_{\mathbb{R}^{2}}\varrho_{\lambda}=1$. Hence, the total energy is
$E_{\lambda}:=\mathcal{E}[\phi_{\lambda}]=K_{\lambda}+U_{\lambda}=\lambda^{2}K_{1}+\lambda^{-2}U_{1},$ (34)
which achieves its minimum at the optimal scaling parameter
$\lambda_{*}:=(U_{1}/K_{1})^{1/4}.$ (35)
Using (34), the corresponding minimal energy is
$E_{\lambda_{*}}=2\sqrt{K_{1}U_{1}}.$ (36)
Gaussian
We now compute the energy $\mathcal{E}[\phi]=\mathcal{E}_{\beta,\gamma,V_{\mathrm{h}}}[\phi]$ analytically for some trial states $\phi$, starting with a radially symmetric, Gaussian state
$\phi(z)=\frac{1}{\sqrt{\pi}}e^{-r^{2}/2},$ (37)
where $r=\lvert z\rvert$. The corresponding density is
$\varrho(z)=\lvert\phi(z)\rvert^{2}=\frac{1}{\pi}e^{-r^{2}},$
satisfying the normalization condition $\int_{\mathbb{R}^{2}}\varrho=1$. Using (12), we get the corresponding magnetic potential
$\mathbf{A}_{\varrho}(z)=\frac{2}{r}\left(\int_{0}^{r}se^{-s^{2}}\,ds\right)\mathbf{e}_{\theta}=\frac{1}{r}(1-e^{-r^{2}})\mathbf{e}_{\theta}.$ (38)
Since $\phi$ is real, $\mathbf{J}_{\phi}=0$. The kinetic energy is
$\int_{\mathbb{R}^{2}}\lvert\nabla\phi\rvert^{2}=2\pi\int_{0}^{\infty}\left\lvert\frac{\partial\phi}{\partial r}\right\rvert^{2}r\,dr=2\int_{0}^{\infty}r^{3}e^{-r^{2}}\,dr=1,$
where we used the standard integral
$\int_{0}^{\infty}x^{2k+1}e^{-ax^{2}}\,dx=\frac{k!}{2a^{k+1}},\qquad k\in\mathbb{N}_{0},\ a>0.$ (39)
The magnetic self-energy is given by
$\int_{\mathbb{R}^{2}}\lvert\mathbf{A}_{\varrho}\rvert^{2}\varrho=2\int_{0}^{\infty}\frac{(1-e^{-r^{2}})^{2}e^{-r^{2}}}{r}\,dr=\log\frac{4}{3},$
which can be derived from the standard integral
$\int_{0}^{\infty}\frac{e^{-ax}-e^{-bx}}{x}\,dx=\log\frac{b}{a},\qquad a,b>0,$ (40)
by expanding the integrand and substituting $t=r^{2}$. The scalar self-interaction energy is
$\int_{\mathbb{R}^{2}}\varrho^{2}=\frac{2}{\pi}\int_{0}^{\infty}re^{-2r^{2}}\,dr=\frac{1}{2\pi},$
and thus
$K_{1}=1+\beta^{2}\log\frac{4}{3}+\frac{\gamma}{2\pi}.$
Finally, the potential energy in the harmonic trap $V(z)=V_{\mathrm{h}}(r)=r^{2}$ is
$U_{1}=\int_{\mathbb{R}^{2}}V_{\mathrm{h}}\varrho=2\int_{0}^{\infty}r^{3}e^{-r^{2}}\,dr=1,$
once again using (39).
Now consider $\phi_{\lambda}$. At the optimal scaling parameter $\lambda=\lambda_{*}$ (see (36)), the total energy is
$\mathcal{E}_{\beta,\gamma,V_{\mathrm{h}}}[\phi_{\lambda_{*}}]=2\sqrt{1+\beta^{2}\log\frac{4}{3}+\frac{\gamma}{2\pi}}.$ (41)
With $\gamma=2\pi\beta\kappa$, this is shown as the dashed blue curve in the plots.
Excited harmonic oscillator eigenstates
The Gaussian is the ground state of the harmonic oscillator. Now, we consider the two angular momentum eigenmodes of the first excited state:
$\phi^{+}(z)=\frac{1}{\sqrt{\pi}}ze^{-\lvert z\rvert^{2}/2},\qquad\phi^{-}(z)=\frac{1}{\sqrt{\pi}}\bar{z}e^{-\lvert z\rvert^{2}/2}.$
In polar coordinates $z=re^{i\theta}$, we have
$\phi^{\pm}(z)=\frac{1}{\sqrt{\pi}}re^{-r^{2}/2}e^{\pm i\theta}.$
The density is the same for both modes:
$\varrho(z)=\lvert\phi^{\pm}(z)\rvert^{2}=\frac{1}{\pi}r^{2}e^{-r^{2}},$
and this is normalized to $\int_{\mathbb{R}^{2}}\varrho=1$. Using (12), the corresponding magnetic potential is $\mathbf{A}_{\varrho}=A(r)\mathbf{e}_{\theta}$, with
$A(r)=\frac{2}{r}\int_{0}^{r}s^{3}e^{-s^{2}}\,ds=\frac{1-(1+r^{2})e^{-r^{2}}}{r}.$
Differentiating $\phi^{\pm}$ yields
$\nabla\phi^{\pm}$ $=\frac{1}{\sqrt{\pi}}e^{\pm i\theta}e^{-r^{2}/2}\left[(1-r^{2})\mathbf{e}_{r}\pm i\mathbf{e}_{\theta}\right]$
$=\frac{\phi_{\pm}}{r}\left[(1-r^{2})\mathbf{e}_{r}\pm i\mathbf{e}_{\theta}\right].$
The currents for the two modes are therefore
$\mathbf{J}_{\pm}:=\mathbf{J}_{\phi^{\pm}}=\pm\frac{\varrho}{r}\mathbf{e}_{\theta}=\pm\frac{1}{\pi}re^{-r^{2}}\mathbf{e}_{\theta}.$
The harmonic oscillator energy contributions are
$\int_{\mathbb{R}^{2}}\lvert\nabla\phi^{\pm}\rvert^{2}=\int_{\mathbb{R}^{2}}V_{\mathrm{h}}\varrho=2,$
which can be verified either by direct computation using (39), or by using that $\phi^{\pm}$ are energy eigenstates of the harmonic oscillator. The scalar self-energy is given by
$\int_{\mathbb{R}^{2}}\varrho^{2}=\frac{2}{\pi}\int_{0}^{\infty}r^{5}e^{-2r^{2}}\,dr=\frac{1}{4\pi}.$
The magnetic self-energy common to both modes is given by
$\int_{\mathbb{R}^{2}}\lvert\mathbf{A}_{\varrho}\rvert^{2}\varrho=2\int_{0}^{\infty}re^{-r^{2}}\left(1-2(1+r^{2})e^{-r^{2}}\right)\,dr=\frac{7}{54},$
again using (39). The contribution due to the current is
$\int_{\mathbb{R}^{2}}\mathbf{J}_{\pm}\cdot\mathbf{A}_{\varrho}=\pm 2\int_{0}^{\infty}re^{-r^{2}}\left(1-(1+r^{2})e^{-r^{2}}\right)\,dr=\pm\frac{1}{4}.$
The total energy is thus
$K_{1}=2+\frac{1}{4\pi}\gamma+\frac{7}{54}\beta^{2}\pm\frac{1}{2}\beta,\qquad U_{1}=2.$
At the optimal scaling parameters $\lambda=\lambda_{\pm}$ (see (36)), we have the minimal energies
$\mathcal{E}_{\beta,\gamma,V_{\mathrm{h}}}\left[\phi^{\pm}_{\lambda_{\pm}}\right]=2\sqrt{4+\frac{\gamma}{2\pi}+\frac{7}{27}\beta^{2}\pm\beta}.$ (42)
Notice however that the optimal scales $\lambda_{\pm}$ differ for the plus and minus modes $\phi^{\pm}$. The negative mode (counter-rotating vortex) gives the lowest energy for $\beta>0$ and is shown as the dashed purple curve in the plots.
.2 Powers of the versiera
Another radially symmetric state is the “versiera” state (20) centered at the origin,
$\phi(z)=\frac{1}{\sqrt{\pi}}\frac{1}{r^{2}+1},$ (43)
which is an NLL state at $\beta=2$. It however has a divergent potential energy. More generally, we can take this state to a positive power $a\geq 1$ (and normalize), yielding the state
$\phi^{a}(z):=\sqrt{\frac{2a-1}{\pi}}\frac{1}{(r^{2}+1)^{a}}.$ (44)
The corresponding density is
$\varrho^{a}(r)=\frac{2a-1}{\pi}\frac{1}{(r^{2}+1)^{2a}},$
which satisfies $\int_{\mathbb{R}^{2}}\varrho^{a}=1$. The corresponding magnetic potential is $\mathbf{A}_{\varrho^{a}}=A(r)\mathbf{e}_{\theta}$, with
$A(r)=\frac{1}{r}\left(1-\frac{1}{(r^{2}+1)^{2a-1}}\right).$
Since $\phi^{a}$ is real, $\mathbf{J}_{\phi^{a}}=0$.
The kinetic energy is
$\int_{\mathbb{R}^{2}}\lvert\nabla\phi^{a}\rvert^{2}=\int_{\mathbb{R}^{2}}\left\lvert\frac{\partial\phi^{a}}{\partial r}\right\rvert^{2}=\frac{2a(2a-1)}{2a+1}.$
The potential energy in a harmonic trap diverges logarithmically for $a=1$, but for $a>1$ we get a finite energy contribution
$\int_{\mathbb{R}^{2}}V_{\mathrm{h}}\varrho^{a}=(2a-1)\int_{0}^{\infty}\frac{r^{2}}{(1+r^{2})^{2a}}\cdot 2r\,dr=\frac{1}{2(a-1)}.$
The scalar self-energy is
$\int_{\mathbb{R}^{2}}(\varrho^{a})^{2}$ $=\frac{1}{\pi}\frac{(2a-1)^{2}}{4a-1}.$
The magnetic self-energy can be expressed using the digamma function
$\psi(z):=\frac{d}{dz}\log\Gamma(z)=\frac{\Gamma^{\prime}(z)}{\Gamma(z)}.$ (45)
Namely, we get
$\int_{\mathbb{R}^{2}}\lvert\mathbf{A}_{\varrho^{a}}\rvert^{2}\varrho^{a}=(2a-1)\vartheta(a),$
where we introduce
$\vartheta(a):=2\psi(4a-1)-\psi(2a)-\psi(6a-2).$ (46)
This can be shown by evaluating the integral
$\int_{\mathbb{R}^{2}}\lvert\mathbf{A}_{\varrho^{a}}\rvert^{2}\varrho^{a}=2(2a-1)\int_{0}^{\infty}\frac{(1-(1+r^{2})^{-(2a-1)})^{2}}{r(1+r^{2})^{2a}}\,dr,$
using a change of variables $t=1+r^{2}$, and the digamma function identity
$\int_{1}^{\infty}\frac{t^{-b}-t^{-a}}{t-1}\,dt=\psi(b)-\psi(a),$ (47)
valid for $a,b\in\mathbb{C}$ with positive real part.
We now consider the rescaled state $\phi^{a}_{\lambda}$. At the optimal scale $\lambda=\lambda_{*}$ (see (36)), we have the total energy
$\mathcal{E}_{\beta,\gamma,V_{\mathrm{h}}}\left[\phi^{a}_{\lambda_{*}}\right]$
$=\sqrt{\frac{2(2a-1)}{a-1}}\sqrt{\frac{2a}{2a+1}+\frac{1}{\pi}\frac{2a-1}{4a-1}\gamma+\vartheta(a)\beta^{2}}.$ (48)
For fixed $(\beta,\gamma)$, this expression can be numerically optimized over $a>1$, yielding an optimal state $\phi^{a_{*}}$. This has been done using Mathematica, and the resulting curve $E=\mathcal{E}[\phi^{a_{*}}]$ for $\gamma=2\pi\beta\kappa$ is the thick yellow line in the plots.
We note that the Gaussian can be re-obtained by e.g. setting $a=2/b$, $\lambda=\sqrt{b/2}$, and taking the limit $b\to 0$. Hence the state $\phi^{a}$ can (for optimal $a>1$) be expected to have a lower energy than the Gaussian state.
We also note that along the critical line $\gamma=-2\pi\beta$, the optimal energy $\mathcal{E}_{\beta,\gamma,V_{\mathrm{h}}}[\phi^{a}_{\lambda_{*}}]$ tends to 0 as $\beta$ tends to the critical value $\beta=2$. This may indeed be derived analytically using the observations
$\psi(z+1)=\psi(z)+\frac{1}{z},\qquad\psi^{\prime}(z+1)=\psi^{\prime}(z)-\frac{1}{z^{2}},$
which follow from the identity $\Gamma(z+1)=z\Gamma(z)$.
Vortex rings
Finally, we consider a symmetric NLL “vortex ring” state of degree (vorticity) n1>0n-1>0, centered at $w=0$:
$\phi(z)=\sqrt{\frac{n}{\pi}}\frac{\bar{z}^{n-1}}{\lvert z\rvert^{2n}+1}.$ (49)
In polar coordinates, $z=re^{i\theta}$, we have
$\phi(z)=\sqrt{\frac{n}{\pi}}\frac{r^{n-1}e^{-i(n-1)\theta}}{r^{2n}+1}.$
The corresponding density
$\varrho(r)=\frac{n}{\pi}\frac{r^{2(n-1)}}{(r^{2n}+1)^{2}}$
is radially symmetric and normalized: $\int_{\mathbb{R}^{2}}\varrho=1$. The magnetic-kinetic self-energy has already been computed in (AtaLunNgu-24, , Prop. 3.35 (proof)) using the supersymmetric factorization method:
$\mathcal{E}_{\beta,\gamma=-2\pi\beta,0}[\phi]$ $=(\beta-2n)^{2}\frac{1}{6}\Gamma\left(1+\frac{1}{n}\right)\Gamma\left(3-\frac{1}{n}\right),$
which, with the scalar self-energy
$\int_{\mathbb{R}^{2}}\varrho^{2}$ $=\frac{n}{6\pi}\Gamma\left(2-\frac{1}{n}\right)\Gamma\left(2+\frac{1}{n}\right),$
yields the total magnetic-kinetic self-energy:
$K_{1}=\mathcal{E}_{\beta,\gamma,0}[\phi]$ $=\mathcal{E}_{\beta,-2\pi\beta,0}[\phi]+(2\pi\beta+\gamma)\int_{\mathbb{R}^{2}}\varrho^{2}.$
The potential energy is
$U_{1}=\int_{\mathbb{R}^{2}}V_{h}\varrho=\Gamma\left(1+\frac{1}{n}\right)\Gamma\left(1-\frac{1}{n}\right)=\frac{\pi/n}{\sin(\pi/n)},$
using the beta function identity
$B(z_{1},z_{2})=\frac{\Gamma(z_{1})\Gamma(z_{2})}{\Gamma(z_{1}+z_{2})}=\int_{0}^{\infty}\frac{t^{z_{1}-1}}{(t+1)^{z_{1}+z_{2}}}\,dt,$
and Euler’s reflection formula,
$\Gamma(1+z)\Gamma(1-z)=\frac{\pi z}{\sin(\pi z)}.$
Hence, at the optimal scaling parameter $\lambda=\lambda_{*}$ (see (36)), the minimal total energy becomes
$\mathcal{E}_{\beta,\gamma,V_{\mathrm{h}}}[\phi_{\lambda_{*}}]$
$=\sqrt{\frac{2}{3}}\sqrt{(\beta-2n)^{2}\left(2-\frac{1}{n}\right)+(2\pi\beta+\gamma)\frac{n+1}{\pi}}$
$\qquad\times\sqrt{1-\frac{1}{n}}\,\frac{\pi/n}{\sin(\pi/n)}.$
For $n\in\{2,3,4,5\}$ (vorticity $1,2,3,4$) and $\gamma=2\pi\beta\kappa$, these are the solid purple, orange, green, respectively blue curves in the plots. Note that indeed $\mathcal{E}\to 0$ linearly as $\beta\to 2n$ and $\kappa=-1$.
Numerical data: Matlab (unconstrained)
Here we tabulate approximate energies of (2) computed in Matlab and optimized using a steepest descent algorithm adopted from CorDubLunRou-19 . As initial states we have started from the states generated in the work CorDubLunRou-19 with $\gamma=0$, and adjusted the parameters $(\beta,\gamma)$ to nearby values in order to generate approximate ground states with different vorticities (the number of vortices listed is approximate in non-symmetric cases). In the stable regime away from criticality, convergence to the desired precision is typically achieved after 50-200 iterations of the algorithm. Due to geometrically and topologically induced metastability, the energy converges into a local minimum, and the global minimum may only be found by comparing a set of reasonable vorticities. Some chosen initial states exhibit sufficient metastability that they can produce estimates for excited states well above the ground levels, by stopping the algorithm after a limited number of iterations during which a stable phase is shown. Thus, all numerically computed energies are only valid as approximate upper bounds to the true ground-state energy, but may also yield indications of higher resonances. At criticality $\kappa=-1$ we have only been able to obtain sufficiently stable and reliable numerics for radially symmetric states. We have also noted some observations, such as approximate linearity, comparison to analytical energies [^1] Note that the constant $c=2\sqrt{\pi}/3$ is a numerical fit approximation to that denoted by $e(1,1)/(2\pi)$ in the earlier work CorDubLunRou-19 ., or cases where we suspect that the state is far from converged. The resulting energies are shown in the plots using solid markers (colors and shapes indicate numbers of visible vortices; see legend in Fig. 4 (bottom)).
$\beta$ $\gamma/\pi$ #vortices $\mathcal{E}[\phi]$ Remark
0 -1.8 0 0.789
-1 0 1.396
0 0 2.000 exact: 2
0 1 4.000 exact: 4
2 0 2.810 $E^{\mathrm{TF}}(\gamma)\approx 1.886$
3.18 0 3.184 BC: 3.1846
15.92 0 5.793 BC: 5.7920
20 0 6.399 $E^{\mathrm{TF}}(\gamma)\approx 5.963$
31.84 0 7.893 BC: 7.8918
79.58 0 12.158 BC: 12.1578
159.16 0 17.024 BC: 17.0236
200 0 19.042 $E^{\mathrm{TF}}(\gamma)\approx 18.856$
2000 0 59.703 $E^{\mathrm{TF}}(\gamma)\approx 59.628$
0.1 -0.2 0 1.900 linear
-0.2 1 3.899 linear?
0 0 2.003
0 1 3.950
0.2 0 2.100 linear
0.2 1 4.001
0.5 -1 0 1.500 linear
-1 1 3.500 linear
0 0 2.071
0 1 3.774
1 0 2.500 linear
1 1 4.030
Table 1: Energies computed in Matlab. BC is comparison to Bao–Cai (BaoCai-12, , Table 3.5) (note factor 2 and possible rounding errors).
$\beta$ $\gamma/\pi$ #vortices $\mathcal{E}[\phi]$ Remark
1 -2 0 1.003 linear?
-2 1 3.013 linear?
-1.8 0 1.204
-1.8 1 3.078
0 0 2.268 $E^{\mathrm{TF}}(2\pi c)\approx 2.050$
0 1 3.607
1 0 2.662
2 0 3.000 linear
2 1 4.120
1.5 -3 0 0.631 unconverged?
-3 1 2.542 unconverged?
0 0 2.560
0 1 3.505
3 0 3.500 linear
3 1 4.265
1.9 -3.8 0 0.188 unconverged?
Table 2: Energies computed in Matlab.
$\beta$ $\gamma/\pi$ #vortices $\mathcal{E}[\phi]$ Remark
2 -4 1 2.024 linear?
-3.6 0 1.081
-3.6 1 2.227
-2 0 2.145
-2 1 2.841
0 0 2.916 $E^{\mathrm{TF}}(4\pi c)\approx 2.899$
0 1 3.474
2 0 3.504
2 1 4.001
4 0 4.000 linear
4 1 4.462
2.5 -5 0 0.736 unconverged?
-5 1 1.553
0 0 3.313
0 1 3.517
5 0 4.500 linear
5 1 4.703
Table 3: Energies computed in Matlab.
$\beta$ $\gamma/\pi$ #vortices $\mathcal{E}[\phi]$ Remark
3 -6 0 1.393
-6 1 1.000 linear
-5.4 0 1.860
-5.4 1 1.506
-5 1 1.761
0 0 3.736
0 1 3.631 $E^{\mathrm{TF}}(6\pi c)\approx 3.550$
6 0 5.000 linear
6 1 4.983
3.5 -7 1 0.504 linear?
Table 4: Energies computed in Matlab.
$\beta$ $\gamma/\pi$ #vortices $\mathcal{E}[\phi]$ Remark
4 -8 0 2.528
-8 1 0.090 unconverged?
-7.2 0 2.829
-7.2 1 1.332
-7.2 2 1.708
-6 1 2.077
-6 2 2.450
-4 1 2.901
-4 2 3.240
-2 1 3.525
-2 2 3.835
0 0 4.631
0 1 4.046 $E^{\mathrm{TF}}(8\pi c)\approx 4.099$
0 2 4.335
0 3 4.536
8 0 6.000 linear
8 1 5.634
8 2 5.878
8 3 6.102
16 1 6.839
16 2 7.062
2000 0 59.840
2000 1 59.890
2000 2 59.970
4.5 -9 1 0.535 linear?
Table 5: Energies computed in Matlab.
$\beta$ $\gamma/\pi$ #vortices $\mathcal{E}[\phi]$ Remark
5 -10 0 3.546
-10 1 1.076 linear?
-9 1 1.850
-9 2 1.679
-9 3 1.805
0 1 4.649
0 2 4.641 $E^{\mathrm{TF}}(10\pi c)\approx 4.583$
0 3 4.826
0 4 5.100
10 1 6.377
10 2 6.402
10 3 6.572
10 4 6.817
Table 6: Energies computed in Matlab.
$\beta$ $\gamma/\pi$ #vortices $\mathcal{E}[\phi]$ Remark
6 -11.8 2 0.715 approx. ring
-10.8 0 4.761
-10.8 1 2.690
-10.8 2 1.708 approx. ring
-10.8 3 1.790
-10.8 4 2.175
-10 2 2.183
-10 3 2.246
-6 1 4.145
-6 2 3.671
-6 3 3.688
-6 4 3.959
0 0 6.510
0 1 5.379
0 2 5.066 $E^{\mathrm{TF}}(12\pi c)\approx 5.021$
0 3 5.078
0 4 5.314
12 0 8.000 linear
12 1 7.183
12 2 6.992
12 3 7.021
12 4 7.220
2000 0 60.011
2000 1 59.985
2000 2 60.025
2000 3 60.081
2000 4 60.145
Table 7: Energies computed in Matlab.
$\beta$ $\gamma/\pi$ #vortices $\mathcal{E}[\phi]$ Remark
7 -13 3 1.561
-12.6 1 3.625
-12.6 2 2.127 2 or 4 vortices (ring)
-12.6 3 1.824
-12.6 4 1.915
0 1 6.188
0 2 5.619
0 3 5.415 $E^{\mathrm{TF}}(14\pi c)\approx 5.423$
0 4 5.518
14 1 8.033
14 2 7.652
14 3 7.522
14 4 7.614
Table 8: Energies computed in Matlab.
$\beta$ $\gamma/\pi$ #vortices $\mathcal{E}[\phi]$ Remark
8 -15 3 1.826 3 or 6 vortices (ring)
-15 4 1.566
-15 6 1.596
-14.4 0 6.721
-14.4 1 4.590
-14.4 3 2.145 3 or 6 vortices (ring)
-14.4 4 1.951 $E^{\mathrm{TF}}(1.6\pi c)\approx 1.833$
-14.4 6 1.993
0 0 8.440
0 1 7.049
0 2 6.274
0 3 5.861
0 4 5.792 $E^{\mathrm{TF}}(16\pi c)\approx 5.797$
16 0 10.000 linear
16 1 8.915
16 2 8.370
16 3 8.086
16 4 8.051 $E^{\mathrm{TF}}(32\pi c)\approx 8.199$
16 5 8.190 not symmetric
Table 9: Energies computed in Matlab.
$\beta$ $\gamma/\pi$ #vortices $\mathcal{E}[\phi]$ Remark
9 -16.2 1 5.582
-16.2 2 3.747 2 or 4 vortices
-16.2 6 2.216 6 or 9 vortices
-16.2 8 2.056 4 or 8 vortices
0 1 7.944
0 2 7.011 2 or 4 vortices
0 4 6.160 $E^{\mathrm{TF}}(18\pi c)\approx 6.149$
0 6 6.298
18 1 9.819
18 2 9.136
18 4 8.544
18 6 8.690
Table 10: Energies computed in Matlab.
$\beta$ $\gamma/\pi$ #vortices $\mathcal{E}[\phi]$ Remark
10 -18 0 8.602 fast collapsing ring
-18 1 6.588
-18 3 3.171 3 or 6 vortices (ring)
-18 6 2.175 $E^{\mathrm{TF}}(2\pi c)\approx 2.050$
-18 8 2.263
-18 9 2.296
-10 6 4.702
-10 8 4.804
0 0 10.392
0 1 8.862
0 2 7.808
0 3 7.040
0 6 6.543 $E^{\mathrm{TF}}(20\pi c)\approx 6.482$
0 8 6.619
10 4 7.967
20 1 10.741
20 3 9.387
20 4 9.095
20 6 9.066 $E^{\mathrm{TF}}(40\pi c)\approx 9.167$
40 4 10.979
80 4 13.959
160 4 18.473
200 4 20.344 $E^{\mathrm{TF}}(220\pi c)\approx 21.498$
2000 4 60.277 $E^{\mathrm{TF}}(2020\pi c)\approx 65.141$
Table 11: Energies computed in Matlab.
$\beta$ $\gamma/\pi$ #vortices $\mathcal{E}[\phi]$ Remark
50 -90 40 4.721 $E^{\mathrm{TF}}(10\pi c)\approx 4.583$
-50 40 10.370 $E^{\mathrm{TF}}(50\pi c)\approx 10.249$
0 40 14.502 $E^{\mathrm{TF}}(100\pi c)\approx 14.494$
100 40 20.228 $E^{\mathrm{TF}}(200\pi c)\approx 20.497$
200 40 24.552 $E^{\mathrm{TF}}(300\pi c)\approx 25.104$
100 -180 88 6.654 $E^{\mathrm{TF}}(20\pi c)\approx 6.482$
-100 88 14.643 $E^{\mathrm{TF}}(100\pi c)\approx 14.494$
0 88 20.499 $E^{\mathrm{TF}}(200\pi c)\approx 20.497$
200 87 28.612 $E^{\mathrm{TF}}(400\pi c)\approx 28.987$
400 88 34.748 $E^{\mathrm{TF}}(600\pi c)\approx 35.502$
Table 12: Energies computed in Matlab.
Numerical data: Mathematica (NLL constraint)
Finally, we use the above analytical solutions to the Wronskian problem for $\beta=2n\in\{4,6,8\}$ to estimate the lowest energy of corresponding NLL states with relaxed radial constraints on the density (a macroscopic, averaged symmetry is still expected due to the radial potential). For $d:=\deg W\in\{2,3,4\}$, we place the vortices in a circle of radius $R\geq 0$: $w_{k}=Re^{i2\pi k/d}$, while for $d=5$ we put one at the origin and the remaining four in a circle; compare (26)–(28). Further, we choose $\eta=0$ due to symmetry, and then numerically optimize the energy over the remaining parameters $\lambda>0$ and $R\geq 0$. This optimization is done in Mathematica by first scanning for reasonable initial values in a rough energy plot and then iteratively rescaling the state as in (34)–(35) in order that the kinetic and potential energies eventually match (convergence towards a fixed point). At given parameter values, the energy is computed by numerical integration of the analytical density of (16) in (10) on a finite domain containing almost all of the mass. The resulting minimal energies are shown in the plots using hollow markers (colors and shapes match the previous vortex numbers; see legend in Fig. 4 (bottom)).
$\beta$ $\deg W$ $\gamma/\pi$ $R$ $\lambda^{-1}$ $\mathcal{E}[\phi_{P,Q}]$ Remark
4 1 -7.2 0 0.8944 1.4049
1 -6 0 1.4142 2.2214
1 -4 0 2.0000 3.1416
1 0 0 2.8284 4.4428
1 8 0 4.0000 6.2832
2 -7.2 0.6054 0.7265 2.7760
2 -6 0.6860 1.0290 4.4693
2 -4 0.8774 1.1965 5.8111
2 0 1.1519 1.3823 7.6549
2 8 1.5034 1.6108 10.3610
6 2 -10.8 0 1.8724 1.7661
2 -6 0 6.2614 3.9487
2 0 0 10.5320 5.5838
2 12 0 17.7194 7.8948
3 -10.8 0.5055 2.0858 2.3043
3 -6 0.8344 3.9167 4.5982
3 0 0.9697 5.6420 6.4626
3 12 1.1428 8.0949 9.1033
4 -10.8 0.8861 0.9747 3.7769 errors
4 -6 1.3621 1.5437 7.8009
4 0 1.5355 2.1498 10.8994
4 12 1.9367 2.4209 14.3487
Table 13: Energies for optimized NLL states.
$\beta$ $\deg W$ $\gamma/\pi$ $R$ $\lambda^{-1}$ $\mathcal{E}[\phi_{P,Q}]$ Remark
8 3 -14.4 0 4.0000 2.2214 errors
3 -14 0 5.0000 2.4360
3 -8 0 20.0010 4.9673
3 0 0 40.0021 7.0247
3 8 0 60.0031 8.6034
3 16 0 80.0042 9.9334
4 -14.4 0.5859 3.4455 2.1411 (a)
4 -14 0.6235 4.0390 2.3895 (b)
4 -8 0.8845 11.3902 4.7762 (c)
4 0 1.0640 19.0020 6.7574 (d)
4 8 1.1502 26.4183 8.2982
4 16 1.2622 32.0333 9.5480
5 -14.4 0.9204 2.5415 2.6768
5 -14 0.8534 2.9129 2.7407
5 -8 1.1760 5.9547 5.4152 errors
5 0 1.4431 8.3299 7.6605
5 8 1.5369 10.4986 9.2824
5 16 1.6036 12.5713 10.7563
Table 14: Energies for optimized NLL states.
Remarks:
Some of the numerical integrations reported errors regardless of precision and may therefore be less reliable.
(a) The energy at the lowest radial vortex ring here ($n=4$) is $\pi/\sqrt{2}\approx 2.2214$.
(b) The energy at the lowest radial vortex ring here ($n=4$) is $\sqrt{5/2}\pi/2\approx 2.4836$.
(c) The energy at the lowest radial vortex ring here ($n=4$) is $\sqrt{5/2}\pi\approx 4.9673$.
(d) The energy at the lowest radial vortex ring here ($n=3$) is $8\sqrt{53}\pi/27\approx 6.7766$.